OpenStudy (anonymous):

I need serious help! what are the foci of the hyperbola with equation 9y^2-16x^2=144

OpenStudy (anonymous):

no one likes these conic section problems, but they are not that hard first of all you need to have an idea of what it looks likek do you know what the picture is?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

ok you need that first lets start by dividing everything by \(144\) to make it look like \[\frac{y^2}{a2}-\frac{x^2}{b^2}=1\]

OpenStudy (anonymous):

you should get \[\frac{x^2}{16}-\frac{y^2}{9}=1\]and now believe it or not we are almost done

OpenStudy (anonymous):

but first you need to know what it looks like 

OpenStudy (anonymous):

any guesses?

OpenStudy (anonymous):

Is it \[(\pm5,0)\]

OpenStudy (anonymous):

No, those a re switched, so it's \[(0,\pm5)\]

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

the picture is the one on the right, since the \(y^2\)term comes first

OpenStudy (anonymous):

so yeah, the foci are \((0,5)\) and \((0,-5)\)